OpenStudy (anonymous):

Please help: 1/2 + 1/8 x 1/5

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } + \frac{ 1 }{ 8} \times \frac{ 1 }{ 5}\]

OpenStudy (anonymous):

The LCD is 40 correct?

OpenStudy (kva1992):

correct

OpenStudy (anonymous):

@kva1992 then do i Multiply 5/40 times 8/40 then plus 20/40 ??? i got it wrong doing it that way

OpenStudy (kva1992):

nope so you do pemdes

OpenStudy (anonymous):

whats does that mean @kva1992

OpenStudy (kva1992):

\[\frac{ 1 }{ 8 } * \frac{ 1 }{ 5 } \]

OpenStudy (kva1992):

its the order of operations

OpenStudy (anonymous):

what about the changing the denominator?

OpenStudy (kva1992):

parentheses exponents multiplication division addition and subtraction and its (pemdas)

OpenStudy (kva1992):

every equation you solve has to go in that order

OpenStudy (kva1992):

so do multiplication first then do the addition it doesnt matter how the question is set up in the beginning

OpenStudy (kva1992):

so it would be like this

OpenStudy (anonymous):

ok,,, great but i am 5/40 x 8/40 plus 20/wrong and its wrong :(

OpenStudy (anonymous):

please show me.... i need the help and for i can use it as future example

OpenStudy (kva1992):

well in multiplication you dont get common denominators

OpenStudy (kva1992):

so like this

OpenStudy (kva1992):